The table below shows whether students from a certain University in Pretoria are right- or left- handed by sex. The total number of students who were interviewed was 100. Their responses were as follows:

    Right handed Left handed Total

Male Female 38 42

12 8

50 50

Total 80 20 100

      3.1.1 What is the probability of selecting a left-handed male? 3.1.2 What is the probability of selecting a right-handed female?

(5) (5)


1
Expert's answer
2021-05-17T00:46:02-0400

Given data is:




3.1.1) "P(Left\ handed\ male)=\dfrac{Left \ Handed \ Male}{Total\ Male}=\dfrac{12}{50}=0.24"


3.1.2) "P(female\ that\ right\ handed)=\dfrac{Female\ of \ right\ handed}{Total\ right\ handed\ students}=\dfrac{42}{80}=0.525"